# EvalTree: Profiling Language Model Weaknesses via Hierarchical Capability Trees

### v1

@@ Abstract
An ideal model evaluation should achieve two goals: identifying where the model fails and providing actionable improvement guidance. Toward these goals for Language Model (LM) evaluations, we formulate the problem of generating a weakness profile, a set of weaknesses expressed in natural language, given an LM’s performance on every individual instance in a benchmark. We introduce a suite of quantitative assessments to compare different weakness profiling methods. We also introduce a weakness profiling method EvalTree. It constructs a capability tree where each node represents a capability described in natural language and is linked to a subset of benchmark instances that specifically evaluate this capability; it then extracts nodes where the LM performs poorly to generate a weakness profile. On the MATH and WildChat benchmarks, we show that EvalTree outperforms baseline weakness profiling methods by identifying weaknesses more precisely and comprehensively. Weakness profiling further enables weakness-guided data collection, and training data collection guided by EvalTree-identified weaknesses improves LM performance more than other data collection strategies. We also show how EvalTree exposes flaws in Chatbot Arena’s human-voter-based evaluation practice. To facilitate future work, we release our code and an interface that allows practitioners to interactively explore the capability trees built by EvalTree.
@@ 1 Introduction
An ideal model evaluation ought to achieve the goals of (1) identifying where the evaluated model fails in a human-interpretable way, and (2) providing actionable guidance to improve the model (Liang et al., 2023; Holtzman et al., 2023; Saxon et al., 2024). However, current model evaluations commonly treat diverse instances in a benchmark uniformly, reducing model performance to a single aggregate metric or coarse-grained, category-level metrics at best. Doing so obscures the reality that a benchmark is heterogeneous, which evaluates diverse capabilities at varying granularities through specific instances, and that model performance can vary significantly across these capabilities. For example, on the MATH benchmark (Hendrycks et al., 2021b), GPT-4o mini (OpenAI, 2024a) achieves an accuracy of 75.1% when calculating combinations and arrangements of elements, but only 49.1% when analyzing geometric relationships using trigonometric principles, as shown in Figure 1(a). As a result, current model evaluations often fail to achieve the two evaluation goals.
@@
Figure 1: (a) EvalTree automatically constructs a capability tree given an LM’s performance on every individual benchmark instance, and then generates a weakness profile by extracting nodes with statistically low performance (weakness profiling). (b) Training data collection guided by weakness profiling effectively improves LM performance, e.g., achieving a 2.5$\times$ accuracy gain on MATH compared to being guided by a generic capability.
@@
Inspired by the preceding observation, we formulate the problem of generating a weakness profile, a set of natural language descriptions of a model’s weaknesses, given the model’s performance on every individual benchmark instance. We focus on profiling Language Model (LM) weaknesses (Figure 1(a)). A weakness (e.g., “analyzing geometric relationships using trigonometric principles”) is a capability where the LM performs poorly in instances that test for this capability. Weakness profiles advance both goals of model evaluation: (1) they provide LM developers and users with an intuitive takeaway to interpret an LM’s heterogeneous performance across diverse capabilities; and (2) they are actionable, e.g., model developers can collect targeted training data to address the identified weaknesses.
@@
In terms of how to profile LM weaknesses, manually analyzing LM performance on all instances is becoming increasingly unrealistic. This is because LM benchmarks are growing in complexity to match the expanding versatility of emerging LMs; moreover, the latest benchmarks such as Chatbot Arena (Chiang et al., 2024) collect real-world human-LM interactions, leading to the emergence of unknown capabilities (Tamkin et al., 2024) tested within a benchmark and thus further complicating manual efforts. Some works thus attempt to automatically profile LM weaknesses by constructing a single-level capability categorization across all benchmark instances and identifying low-performing categories (Murahari et al., 2024; Moayeri et al., 2024); however, fixed-granularity categorizations could be either too broad to provide precise diagnoses or too specific to retain high-level interpretability. More critically, while some methods, including those mentioned above, have been qualitatively shown to identify LM weaknesses, there is no existing study to compare them quantitatively.
@@
Our experiments show that EvalTree advances both evaluation goals via weakness profiling: (1) EvalTree profiles LM weaknesses more precisely and comprehensively than existing methods on the MATH and WildChat (Zhao et al., 2024a) benchmarks; (2) synthetic data generation guided by EvalTree-identified weaknesses effectively improves LM performance, e.g., achieving a 2.5$\times$ accuracy gain on MATH compared to being guided by a generic capability (Figure 1(b)). Furthermore, we show how EvalTree uncovers abnormal LM rankings in Chatbot Arena, exposing flaws in its human-voter-based evaluation practice. We also provide an interface that lets practitioners interactively explore capability trees to facilitate future work. Finally, we discuss future directions, including improving capability tree construction and leveraging capability trees for various potential applications.
@@ 1.1 Related Work
Some prior work explores how to identify LM weaknesses by constructing custom instance sets to specifically highlight underperforming areas (Ribeiro & Lundberg, 2022; Gao et al., 2023; Li et al., 2024). In contrast, we operate entirely on existing benchmarks and emphasize interpretability. In terms of methodology, while EvalTree automatically constructs a tree to organize instances in a dataset, a small number of datasets are released with similar hierarchical structures defined by their creators. For example, several datasets provide shallow trees, e.g., a two-layer taxonomy (Wang et al., 2022; Bai et al., 2024; Zhong et al., 2024a); some adopt existing trees to guide data collection, such as ImageNet (Deng et al., 2009) using WordNet (Miller, 1994) and iNat2017 (Horn et al., 2018) using a biological taxonomy. Some prior work also studies structured capability categorization, the essential idea behind EvalTree; e.g., QualEval (Murahari et al., 2024) and Skill-Slices (Moayeri et al., 2024) propose LM-based pipelines to automatically categorize benchmark instances into capability groups, providing single-level capability categorization structures. Most related to our work, Wang et al. (2023); Zhong et al. (2024b) suggest recursively clustering instances in a dataset to construct trees, and Anthropic’s internal system Clio (Tamkin et al., 2024) employs Claude to build trees of human-LM conversations based on specific attributes or characteristics (e.g., topic). However, these techniques either incur prohibitively high LM usage costs or do not release key implementation details and source code, making them difficult to use. Most importantly, those works do not demonstrate how methods based on their trees can be quantitatively compared with other methods on concrete problems.
@@ 2.1 Definition and Desiderata
    Low-performance identification (precision): The LM should exhibit low performance on instances (sampled from $\mathcal{D}$) testing for each identified weakness.
@@
A weakness profiling method takes as input an LM’s evaluation result on a given benchmark of size $N$ sampled from the data distribution $\mathcal{D}$, represented as a vector $g\in\mathbb{R}^{N}$, where each $g_{i}$ denotes the performance metric achieved by the LM on the $i$-th instance. We refer to this instance set as the profiling set. Since “weakness” is inherently a relative concept, a weakness profiling method should also include a user-tunable hyperparameter $\tau$ to control strictness, where a higher $\tau$ results in weaknesses being identified at higher performance levels. For example, one might set $\tau$ higher to focus on general areas for improvement, while another might adjust $\tau$ lower to find the LM’s extreme failures. When referring to a specific method in context, we denote $W_{\tau}$ as the weakness profile generated with a given $\tau$.
@@ 2.2 Assessment for Comparing Weakness Profiles
We assume the existence of a test set sampled from the data distribution $\mathcal{D}$. We denote the LM’s evaluation result vector on this test set as $f$, analogous to $g$ defined above for the profiling set. We also define the LM’s performance metric over a set of instance indices $S$ as $F(S)=\sum_{x\in S}f_{x}/|S|$, assuming that the performance metric can be averaged; for example, each $f_{i}$ might be a binary value (0/1) indicating whether the LM correctly solved the $i$-th instance, in which case $F(S)$ is the accuracy of the LM on the set $S$. Furthermore, given a capability description $c\in\mathcal{C}$, we call an instance that tests for this capability an associated instance of $c$, with the index set of all associated instances in the test set denoted as $A(c)$. In our experiments, we prompt an LM to determine whether a given instance is an associated instance of a capability $c$ to get $A(c)$, with further details in Appendix E.1.
@@
Low-Performance Identification Assessment. To measure desideratum 1, i.e., low-performance identification, we measure how low $\sum_{w_{i}\in W}F(A(w_{i}))/|W|$ can be, i.e., the (average) performance metric on instances that test for an identified weakness $w_{i}$. Denoting $S=\bigcup_{w_{i}\in W}A(w_{i})$, we also compare how low $F(S)$ can be, i.e., the performance metric on all instances that test for at least one weakness in $W$. In the two cases, a lower value indicates weaker performance in the identified weaknesses, which can better satisfy desideratum 1.
Ground-Truth Weakness Assessment. To measure all three desiderata, inspired by Zhong et al. (2023), we generate a synthetic evaluation result for a “hypothetical” LM’s performance on the profiling set. We use synthetic evaluation results rather than evaluation results of real LMs because desideratum 2, i.e., comprehensive coverage, cannot be reliably measured without prior knowledge of the LM’s true weaknesses, which is exactly the problem we are trying to solve; by generating a synthetic evaluation result, we can control the ground-truth weaknesses, allowing for a rigorous assessment. We start with a predefined ground-truth weakness profile $W^{*}=\{w_{1}^{*},w_{2}^{*},\ldots,w_{M^{*}}^{*}\}$. Then, we sample each $g_{i}$, i.e., the LM’s performance metric on the $i$-th benchmark instance (in the profiling set), ensuring that instances associated with weaknesses in $W^{*}$ exhibit systematically lower performance than other instances; specifically, we independently sample each $g_{i}$ such that instances associated with weaknesses in $W^{*}$ tend to have lower values of $g_{i}$ than other instances. Finally, to assess a weakness profile $W$, we measure its alignment with $W^{*}$ based on the overlap of associated instances in the test set; we restrict $|W|$ to values that are not significantly larger than $|W^{*}|$, preventing methods from inflating scores by generating overly specific descriptions that increase $|W|$, which would violate desideratum 3, i.e., appropriate granularity.
Extrinsic Assessment: Weakness-Guided Training Data Collection. We examine the effectiveness of a weakness profile in supporting the second evaluation goal of improving the evaluated LM. In real-world scenarios, LM developers collect additional finetuning data and perform continual training to further improve an LM. A common strategy is to collect data guided by a generic capability such as “mathematical reasoning”. We hypothesize that a weakness-guided strategy, wherein a weakness profile for the LM is used as actionable guidance for targeted data collection, may be more effective by directly addressing where the LM fails. For a controlled comparison, we collect data by synthetic data generation and compare LMs trained on data generated with different weakness profiles.
@@ 3.2 Generating a Weakness Profile from the Capability Tree
EvalTree generates an LM weakness profile by extracting nodes where the LM’s performance metric is significantly below a user-tunable threshold $\tau$; for clarity, we consider the specific case of correctness-based accuracy being the metric. The extraction algorithm traverses the capability tree from the root to the leaves (for further details, see Appendix B):
@@
    Statistical Test. At each visited node, we perform a binomial test to determine whether its accuracy is significantly lower than $\tau$. The test uses the number of linked instances as the total sample size and the number of correctly solved instances as the count of successes. We apply the same test to the node’s direct children.
@@
Finally, the nodes extracted from running the algorithm are non-overlapping, i.e., no instance (leaf node) is linked to more than one extracted node. The final weakness profile consists of the capability descriptions of the extracted nodes. By adjusting the meaning of “count of successes” in the statistical test, this algorithm also supports various metrics and can identify strengths (performance above a threshold). Note that setting a significance threshold of $1-\alpha$ at each node’s statistical test does not guarantee an overall $1-\alpha$ confidence level across all tests conducted at multiple nodes since we do not correct the tests for multiple comparisons; we address this by incorporating any adjustments into the choice of $\tau$.
@@ 4 Baseline Methods for Profiling LM Weaknesses
TextDiff. TextDiff (Zhong et al., 2022) is an LM-based method that automatically describes differences between two text distributions in natural language. While not originally designed for weakness profiling, prior work has used it to describe distributional differences between two instance sets. We adapt this method by comparing instances where the evaluated LM fails versus succeeds, using the described differences to identify its weaknesses. Specifically, we randomly sample two sets of instances: those where the evaluation result indicates that the evaluated LM has failed, and those where it has succeeded. We then prompt a diagnostic LM using the sampled instances to output a predefined number of potential weaknesses that might cause the evaluated LM to struggle. We compute the performance on the associated instances in the profiling set (Section 2.2) for each potential weakness and select those with the lowest performance metrics as the weakness profile. Note that this step actually gives TextDiff an unfair advantage over other methods in our experiments, as it uses the same implementation used by the method assessment to determine associated instances; however, a method should not have access to this information in principle, such as which LM is used or what prompt is used for method assessment.
QualEval. QualEval (Murahari et al., 2024) uses an automatic LM-based pipeline to derive a predefined number of capabilities (e.g., 20) described in natural language from all benchmark instances. The method then applies a linear programming algorithm to assign each benchmark instance to some of the derived capabilities. Finally, it outputs a single-level capability categorization structure. We compute the performance metric on all instances assigned to each capability and identify a set of weaknesses as the weakness profile by selecting capabilities with the lowest performance metrics.

### v2

@@ Abstract
An ideal model evaluation should achieve two goals: identifying where the model fails and providing actionable improvement guidance. Toward these goals for language model (LM) evaluations, we formulate the problem of generating a weakness profile, a set of weaknesses expressed in natural language, given an LM’s performance on every individual instance in a benchmark. We introduce a suite of quantitative assessments to compare different weakness profiling methods. We also introduce a weakness profiling method EvalTree. EvalTree constructs a capability tree where each node represents a capability described in natural language and is linked to a subset of benchmark instances that specifically evaluate this capability; it then extracts nodes where the LM performs poorly to generate a weakness profile. On the MATH and WildChat benchmarks, we show that EvalTree outperforms baseline weakness profiling methods by identifying weaknesses more precisely and comprehensively. Weakness profiling further enables weakness-guided data collection, and training data collection guided by EvalTree-identified weaknesses improves LM performance more than other data collection strategies. We also show how EvalTree exposes flaws in Chatbot Arena’s human-voter-based evaluation practice. To facilitate future work, we provide an interface that allows practitioners to interactively explore the capability trees built by EvalTree.
@@ 1 Introduction
An ideal model evaluation ought to achieve the goals of (1) identifying where the evaluated model fails in a human-interpretable way, and (2) providing actionable guidance to improve the model (Liang et al., 2023; Holtzman et al., 2023; Gu et al., 2024; Saxon et al., 2024). However, current model evaluations commonly treat diverse instances in a benchmark uniformly, reducing model performance to a single aggregate metric or coarse-grained, category-level metrics at best (Raunak et al., 2022). Doing so obscures the reality that a benchmark is heterogeneous, which evaluates diverse capabilities at varying granularities, and that model performance can vary significantly across these capabilities. For example, on the MATH benchmark (Hendrycks et al., 2021b), GPT-4o mini (OpenAI, 2024a) achieves an accuracy of 75.1% when calculating combinations and arrangements of elements, but only 49.1% when analyzing geometric relationships using trigonometric principles, as shown in Figure 1(a). As a result, current model evaluations often fail to achieve the two goals.
@@
Figure 1: (a) EvalTree automatically constructs a capability tree given an LM’s performance on every individual benchmark instance, and then generates a weakness profile by extracting tree nodes with statistically low performance. (b) Training data collection guided by weakness profiling effectively improves LM performance, e.g., achieving an accuracy gain that is 2.5$\times$ larger than that obtained when being guided by a generic capability.
@@
Inspired by the preceding observation, we formulate the problem of generating a weakness profile, a set of natural language descriptions of a model’s weaknesses, given the model’s performance on every individual benchmark instance. We focus on profiling language model (LM) weaknesses (Figure 1(a)). A weakness (e.g., “analyzing geometric relationships using trigonometric principles”) is a capability where the LM performs poorly on instances that test for this capability. Weakness profiles advance both goals of model evaluation: (1) they provide practitioners with an intuitive takeaway to interpret where an LM fails, based on its heterogeneous performance across diverse capabilities; and (2) they are actionable, e.g., model developers can collect targeted training data to address the identified weaknesses.
@@
In terms of how to profile LM weaknesses, manually analyzing LM performance on all instances is becoming increasingly unrealistic. Some works thus attempt to automatically profile LM weaknesses by constructing a single-level capability categorization across all benchmark instances and identifying low-performing categories (Murahari et al., 2024; Moayeri et al., 2024); however, fixed-granularity categorizations could be either too broad to provide precise diagnoses or too specific to retain high-level interpretability. More critically, while some methods, including those mentioned above, have been qualitatively shown to identify LM weaknesses, there is no existing study to compare them quantitatively.
@@
Our experiments show that EvalTree advances both evaluation goals via weakness profiling: (1) EvalTree profiles LM weaknesses more precisely and comprehensively than existing methods on the MATH and WildChat (Zhao et al., 2024a) benchmarks; (2) synthetic data generation guided by EvalTree-identified weaknesses effectively improves LM performance, e.g., achieving an accuracy gain that is 2.5$\times$ larger than that obtained when being guided by a generic capability (Figure 1(b)). Furthermore, we show how EvalTree uncovers abnormal LM rankings in Chatbot Arena, exposing flaws in its human-voter-based evaluation practice. We also provide an interface that lets practitioners interactively explore capability trees to facilitate future work. Finally, we discuss future directions, including improving capability trees and leveraging capability trees for potential applications.
@@ 1.1 Related Work
Structured Categorization. Structured categorization of benchmark instances is the essential idea behind EvalTree. Murahari et al. (2024); Moayeri et al. (2024) automatically categorize benchmark instances into capability groups, providing single-level capability categorization structures. A small number of datasets are released with hierarchical structures defined by their creators. For example, some provide shallow trees, e.g., a two-layer taxonomy (Wang et al., 2022; Bai et al., 2024; Zhong et al., 2024a); some adopt existing trees to guide data collection, such as ImageNet (Deng et al., 2009) using WordNet (Miller, 1994) and iNat2017 (Horn et al., 2018) using a biological taxonomy. Most related to our work, Wang et al. (2023); Zhong et al. (2024b) recursively cluster instances in a dataset to construct trees, and Anthropic’s internal system Clio (Tamkin et al., 2024) employs Claude 3.5 (Anthropic, 2024) to build trees of human-LM conversations based on specific attributes or characteristics (e.g., topic). However, these techniques either incur prohibitively high LM usage costs or do not release key implementation details and source code, making them difficult to use.
Automatic Weakness Identification. Manually analyzing LM performance on instances in a benchmark for weakness profiling is becoming increasingly unrealistic. This is because LM benchmarks are growing in complexity to match the expanding versatility of emerging LMs; moreover, some datasets (e.g., WildChat (Zhao et al., 2024a)) collect real-world human-LM interactions, leading to the emergence of capabilities (tested within the benchmark) that are not foreseeable even by their creators in advance, further complicating manual efforts. Some works thus attempt to automatically profile LM weaknesses by using LMs to analyze evaluation results (Zhong et al., 2022) or by identifying low-performing categories from a single-level capability categorization (Murahari et al., 2024; Moayeri et al., 2024). Among these works, we are the first to formulate the problem of weakness profiling with quantitative assessments. Targeting similar goals, some works identify interpretable weaknesses (Eyuboglu et al., 2022; Hua et al., 2023), but assume closed output spaces, making them unsuitable for open-ended tasks; others (Wu et al., 2019; Ribeiro et al., 2020) propose interactive tools based on predefined failure modes, whereas we aim for fully automated profiling without such assumptions. Separately, while weakness profiling operates entirely on existing benchmarks and emphasizes interpretability, some prior work explores identifying model weaknesses by constructing custom instance sets to highlight underperforming areas (Ribeiro & Lundberg, 2022; Gao et al., 2023; Li et al., 2024; Wang et al., 2025).
@@ 2.1 Definition and Desiderata
    Low-performance identification (precision): The LM should exhibit low performance on instances (sampled from $\mathcal{D}$) testing for each identified weakness $w_{i}\in W$.
@@
A weakness profiling method takes as input an LM’s evaluation result on a given benchmark of size $N$ sampled from the data distribution $\mathcal{D}$, represented as a vector $g\in\mathbb{R}^{N}$, where each $g_{i}$ denotes the performance metric achieved by the LM on the $i$-th instance. We refer to this instance set as the profiling set. Since “weakness” is inherently a relative concept, a weakness profiling method should also include a user-tunable hyperparameter $\tau$ to control strictness; for example, increasing $\tau$ makes weakness identification less strict, allowing capabilities with relatively higher performance to be identified, whereas decreasing $\tau$ makes it more strict, restricting identification to the LM’s most severe failures. When referring to a specific method in context, we denote $W_{\tau}$ as the weakness profile generated with a given $\tau$.
@@ 2.2 Assessment for Comparing Weakness Profiles
We assume the existence of a test set sampled from the data distribution $\mathcal{D}$. Furthermore, given a capability description $c\in\mathcal{C}$, we call an instance that tests for this capability an associated instance of $c$, with the index set of all associated instances in the test set denoted as $A(c)$. In our experiments, we prompt an LM to determine whether a given instance is an associated instance of a capability $c$ to get $A(c)$, with further details in Appendix E.1.
@@
Low-Performance Identification Assessment. We denote the LM’s evaluation result vector on the test set as $f$, analogous to $g$ defined above for the profiling set. We also define the LM’s performance metric over a set of instance indices $S$ as $F(S)=\sum_{x\in S}f_{x}/|S|$, assuming that the performance metric can be averaged; for example, each $f_{i}$ might be a binary value (0/1) indicating whether the LM correctly solved the $i$-th instance, in which case $F(S)$ is the accuracy of the LM on the set $S$. To measure desideratum 1, i.e., low-performance identification, we examine how low the average performance across identified weaknesses can be, computed as $\sum_{w_{i}\in W}F(A(w_{i}))/|W|$. Denoting $S=\bigcup_{w_{i}\in W}A(w_{i})$, we also compare how low $F(S)$ can be, i.e., the performance metric on all instances that test for at least one identified weakness in $W$. In the two comparisons, a lower metric value indicates weaker performance on the identified weaknesses, which can better satisfy desideratum 1.
Ground-Truth Weakness Assessment. To measure all three desiderata, inspired by Zhong et al. (2023), we generate a synthetic evaluation result for a “hypothetical” LM’s performance on the profiling set. We use synthetic evaluation results rather than evaluation results of real LMs because desideratum 2, i.e., comprehensive coverage, cannot be reliably measured without prior knowledge of the LM’s true weaknesses, which is exactly the problem we are trying to solve. By generating a synthetic evaluation result, we can control the ground-truth weaknesses and thus have such prior knowledge, allowing for a rigorous assessment. We start with a predefined ground-truth weakness profile $W^{*}=\{w_{1}^{*},w_{2}^{*},\ldots,w_{M^{*}}^{*}\}$. Then, we independently sample each $g_{i}$ such that instances associated with weaknesses in $W^{*}$ have systematically lower values of $g_{i}$ than others. Finally, to assess a weakness profile $W$, we measure its alignment with the ground-truth profile $W^{*}$ based on the overlap of associated instances in the test set; we restrict $|W|$ to values that are not significantly larger than $|W^{*}|$, preventing methods from inflating scores by generating overly specific descriptions that increase $|W|$, which would violate desideratum 3, i.e., appropriate granularity.
Extrinsic Assessment: Weakness-Guided Training Data Collection. We examine the effectiveness of a weakness profile in supporting the second evaluation goal of improving the evaluated LM. In the real world, LM developers collect additional training data and perform finetuning to further improve an LM. A common strategy is to collect data guided by a generic capability such as “mathematical reasoning”. We hypothesize that a weakness-guided strategy, wherein a weakness profile for the LM serves as actionable guidance for targeted data collection, may be more effective by directly addressing where the LM fails. For a controlled comparison, we collect data by synthetic data generation and compare LMs trained on data generated under the guidance of different weakness profiles.
@@ 3.2 Generating a Weakness Profile from the Capability Tree
EvalTree generates an LM weakness profile by extracting nodes where the LM’s performance metric is significantly below a user-tunable threshold $\tau$; for clarity, we consider the specific case of correctness-based accuracy being the metric. The extraction algorithm traverses the capability tree from the root to the leaves (see Appendix B for details):
@@
    Statistical Test. At each visited node, we perform a binomial test to determine whether its accuracy is significantly lower than $\tau$. The test uses the number of linked instances as the total sample size and the number of correctly solved instances as the count of successes. We apply the same test to the node’s direct children [^1] Note that setting a significance level of $\alpha$ for each node’s statistical test does not guarantee an overall $1-\alpha$ confidence level across all tests, as they are not corrected for multiple comparisons..
@@
Finally, the nodes extracted from running the algorithm are non-overlapping, i.e., no instance (leaf node) is linked to more than one extracted node. The final weakness profile consists of the capability descriptions of the extracted nodes. By adjusting the meaning of “count of successes” in the statistical test, this algorithm also supports various metrics (e.g., accuracy and win-rate) and can identify strengths (performance above a threshold).
@@ 4 Baseline Methods for Profiling LM Weaknesses
TextDiff (Zhong et al., 2022) is an LM-based method that automatically describes differences between two text distributions in natural language. While not originally designed for weakness profiling, prior work has used it to describe distributional differences between two instance sets. We adapt this method by comparing instances where the evaluated LM fails versus succeeds, using the described differences to identify its weaknesses. Specifically, we randomly sample two sets of instances: those where the evaluation result indicates that the evaluated LM has failed, and those where it has succeeded. We then prompt a diagnostic LM using the sampled instances to output a predefined number of potential weaknesses that might cause the evaluated LM to struggle. We compute the evaluated LM’s performance on the associated instances in the profiling set (Section 2.2) for each potential weakness and select those with the lowest performance metrics as the weakness profile. Note that this step actually gives TextDiff an unfair advantage over other methods in our experiments, as it uses the identical implementation used by the method assessment to determine associated instances; however, a method should not have access to this information in principle, such as which LM is used or what prompt is used for method assessment.
QualEval (Murahari et al., 2024) uses an automatic LM-based pipeline to derive a predefined number of capabilities (e.g., 20) described in natural language from all benchmark instances. The method then applies a linear programming algorithm to assign each benchmark instance to some of the derived capabilities. Finally, it outputs a single-level capability categorization structure. We compute the evaluated LM’s performance metric on all instances (in the profiling set) assigned to each capability and identify a set of weaknesses as the weakness profile by selecting capabilities with the lowest performance metrics.